In a triangle ABC,a=4,b=3,A=60.Then c is the root of the equation
  1. c^2-3c-7=0
  2. c^2+3c+7=0
  3. c^2-3c+7=0
  4. c^2+3c-7=0

Grade:12

1 Answers

moumi roy
7 years ago
answer is (a).
using cos rule:a^2=b^2+c^2-2b*c*cosA
putting values:16=9+c^2-2*3*c*(.5)
we get:c^2-3c-7=0